# Defending Against Malicious Behaviors in Federated Learning with Blockchain

### v1

@@ 1 Introduction
Though integrating blockchain with existing FL systems can partially solve the threat to the central server, FL systems are still vulnerable to client-side malicious attacks [14]. In this work, we define malicious behaviors as actions that intentionally decrease the learning performance (e.g. missing accuracy and convergence) of the global model. The attackers can sabotage the FL systems via attacks such as data poisoning [15] or model poisoning [14]. This work focuses on defending against client-side malicious attacks.
@@
We propose a generic framework that can integrate an FL system with a blockchain system and can defend against malicious attacks. The proposed defense mechanism is motivated by proof-of-stake (PoS) [16], a consensus mechanism in blockchain, and The Resistance [17], a role-playing board game. PoS has an incentive mechanism that encourages honest behaviors by rewarding it and punishes dishonest behaviors via slashing. The Resistance, on the other hand, has two mismatched competing parties, where the party with a larger size is denoted as the resistance force and the other party is denoted as the spies. In The Resistance, there is a voting mechanism where, in each round, each player conducts independent reasoning and votes for a player, and the player with the highest votes will be deemed as a “spy” and kicked out of the game. The goal of the resistance force is to vote out all the spies while the spies aim to impersonate the resistance force and survive until the end. Based on these two concepts, this work proposes a novel majority-voting mechanism for global aggregation where each participating client independently validates the quality of aggregated local updates and votes for acceptance of the global update. The aggregation mechanism is stake-based where participating clients stake assets [^1] In practice, the staked assets can be linked with cryptocurrency or real currency to increase the financial cost of malicious attacks. or tokens (a quantitative measurement of the asset, which can be used to indicate the trustworthiness of the client in our system) for their own actions. There are two types of actions, proposing (uploading local updates) and voting. If the majority vote is to accept the global aggregation, a proposer will be refunded with its staked tokens and a voter who votes for acceptance will not only be refunded but also be rewarded with the staked tokens from the voters who vote for rejection, and vice versa. The overall procedure of the stake-based aggregation mechanism is illustrated in Fig. 1.
@@
We evaluate the proposed framework on a practical financial problem, namely loan default prediction. We simulate the FL and blockchain environment for the Lending Club Kaggle challenge dataset to conduct experiments in a controllable setting and to provide insights into the problem of interest. We empirically show that an FL system can maintain robust performance under malicious attacks by introducing the proposed stake-based aggregation mechanism.
@@
We formulate the problem of decentralized federated learning with blockchain in the presence of malicious attacks.
We propose a stake-based aggregation mechanism for federated learning systems that can defend against malicious attacks.
We evaluate the robustness of the proposed framework in a simulated environment and provide initial empirical insights into the problem.
@@ 2 Related Work
2.1 Federated Learning
The concept of FL comes from the necessity of on-device training, where the training data have to remain on the device [6]. The clients of FL are distributed at different physical locations that are connected to the internet, which exposes a few security risks compared with distributed learning. First, as the local dataset belongs to the client, FL has to take the users’ privacy into consideration. This can be addressed by integrating privacy-preserving techniques into FL, such as differential privacy [18]. Second, FL can be manipulated via internet access, e.g. missing the central server can be compromised by a third party. Third, a new client could participate in the federated training at any time if it meets the required criteria. This means that clients with malicious intentions can also join the federated systems while meeting the initial criteria. This work focuses on the third risk as the second risk is mitigated by replacing the central server with a blockchain. Traditional FL methods can only detect and defend the malicious clients via assigning small weights to malicious clients during global aggregation based on the divergence of learned parameters [19] or using unsupervised learning methods (e.g. missing anomaly detection [20] and clustering [21]) to get rid of malicious clients. However, none of these methods tackles the second challenge and these methods do not consider the third challenge. The proposed framework utilizes the blockchain to ensure the security of global aggregation and defends the client-side malicious behaviors with a novel majority-voting mechanism.
@@ 2.2 Blockchain
Blockchains refer to distributed ledgers that operate on a global peer-to-peer (P2P) network, as exemplified by popular cryptocurrencies such as Bitcoin [22] and Ethereum [9]. One of the defining characteristics of blockchain technology is the ability for users to freely join or leave the network, without a central authority in place to ensure common agreement on the distributed ledgers. Instead, users rely on consensus protocols [16, 23], such as proof-of-work (PoW) or PoS, to achieve agreement in a distributed setting.
As shown in Fig. 2, in a blockchain system, a transaction typically involves a sender who wishes to transfer a digital asset, such as a cryptocurrency, to a recipient. The sender initiates the transaction by creating a digital signature that includes the transaction details and the sender’s private key, which is used to verify the sender’s identity and authorize the transfer. The transaction is then broadcasted over a P2P network to miners, who are participants in the network responsible for verifying and adding new blocks of transactions to the blockchain. Miners validate and confirm the transaction using consensus protocols, to ensure that the transaction is legitimate and not a duplicate or fraudulent transaction. Once confirmed, the transaction is added to a block, which is then cryptographically linked to the previous block using hash functions [24], forming a chain of blocks (i.e. missing, blockchain). The block is then propagated to all the participants in the network, creating a decentralized, immutable record of the transaction. Finally, the recipient can access the digital asset by using their private key to authenticate their identity and claim ownership of the asset. The use of cryptography and consensus protocols ensures the security, transparency, and decentralization of the transaction process, making blockchain technology a promising solution for a variety of applications beyond cryptocurrency, e.g., insurance [25], healthcare [11], supply chain management [12], energy [26], and Internet of Things (IoT) [27].
Another key feature of blockchain technology is the use of smart contracts [9], which are quasi-Turing-complete programs that can be executed within a virtual machine. When a transaction is initiated, a smart contract is typically used to encode the terms and conditions of the transaction, such as the amount, currency, and time of transfer. The smart contract is then stored on the blockchain network and executed automatically when the predefined conditions are met. The execution of the smart contract verifies the transaction, ensuring that it meets the agreed-upon terms and conditions, and then automatically transfers the digital asset or currency to the recipient. Smart contracts can be leveraged to build a wide range of decentralized applications (DApps), such as decentralized finance (DeFi) services [28].
@@ 2.3 Federated Learning with Blockchain
Traditional FL faces challenges [29], such as privacy and security concerns, unreliable communication, and difficulty in reaching a consensus among the parties. Blockchain, on the other hand, provides a decentralized, secure, and transparent platform for data storage and sharing. This makes the use of blockchain for FL a promising direction to potentially address privacy and security concerns by allowing parties to keep their data private while still contributing to the training process. Additionally, blockchain can provide a secure communication channel for FL participants and ensure the integrity of the FL process.
Current blockchain-based FL designs [30, 8, 31, 32] have been broadly used in diverse fields, including mobile edge computing [33], IoT [34] and distributed machine learning [35]. Despite the potential benefits of combining FL with blockchain, several challenges remain. For instance, FL systems are still vulnerable to client-side malicious attacks [14] and lack incentive-compatible mechanisms to motivate FL participants to behave honestly during the training process. Multiple reputation-based incentive mechanisms [36, 37] have recently been proposed to encourage participants and enhance model accuracy in blockchain-based FL. However, it remains unclear how to effectively utilize the blockchain infrastructure and leverage its inherent incentive mechanism (i.e., cryptocurrencies) to incentivize trustworthy FL behaviors and penalize malicious clients.
@@ 3 Problem Formulation
This section introduces the problem of interest, the definition of the malicious behaviors considered, and the underlying assumptions in this work.
@@ 3.3 Assumptions
A4: There is no collusion among malicious clients. That is to say, each malicious client acts independently. In the application scenarios of this work (e.g. missing Sec. 5), there is a minimal bar for a client to participate in the system, e.g. missing basic qualifications or industry standards for clinical or financial institutes. Meanwhile, it is difficult to compromise multiple clients with independent cybersecurity systems simultaneously. Thus, we deem that it is almost impossible to launch large-scale multi-agent attacks in the application scenarios of interest. As an exploratory study, this work considers the single-agent scenario. (See Sec. 4.5 for the discussion on a multi-agent scenario.)
A5: There is no capacity constraint on the hardware, including computing, communication, and storage, allowing us to solely focus on the algorithmic side of the problem.
A6: The underlying blockchain of the FL system of interest is running securely with a consensus protocol that ensures the validity and integrity of transactions and blocks. While the security of the blockchain is crucial for the overall security of the FL system, addressing the malicious miners falls outside the scope of this study.
@@ 4 Method
Figure 3: A round-based training process. In the initial state (indexed as ①), both honest (black) and malicious (red) clients exist in an FL system. In the final state (indexed as ⑥), all malicious clients are expected to be removed from the system. To reach the final state from the initial state, multiple rounds of training are required. Here are two possible scenarios, the proposed aggregation is either approved (the upper branch) or denied (the lower branch) by the voters. In each round (within the dotted orange line), a subset of clients are randomly selected as proposers, and another subset of clients are randomly selected as voters. The proposers and voters interact with the blockchain following the order of orange arrows (from ② to ⑤).
@@ 4.2 Local Validation
In contrast to standard FL algorithms, the aggregated global model is not recorded in a block directly. Instead, $\tilde{\theta}_{0}^{t}$, a copy of $\theta_{0}^{t}$ is downloaded by a randomly selected set of clients, denoted as voters, $\mathcal{K}_{v}^{t}$. A voter $k$ runs a local inference with $\tilde{\theta}_{0}^{t}$ on its local test set and outputs a local validation score. The local validation score $s^{t}_{k}$ is a scalar, which can be linked with common metrics of ML tasks [^2] For example, common evaluation metrics include accuracy for classification, mean Intersection over Union (mIOU) for semantic segmentation, and mean average precision (mAP) for object detection.. If $s^{t}_{k}$ is not lower than a threshold, the voter votes for accepting this aggregated model; otherwise, the voter votes against it. The threshold can be based on a validation score $s^{t-1}_{k}$ acquired in the previous round. In the training of ML tasks, the scores can be volatile due to the characteristics of the tasks. Thus, a hyperparameter $\epsilon\in(0,1)$ is introduced to control the tolerance of performance decrease in a single round. Mathematically, Voter $k$ has the following score.
@@
It is worth mentioning that it is almost impossible for the attackers to manipulate scores by fooling all the randomly selected voters (e.g. missing via adversarial attacks [40]). According to A5, the majority of voters are honest. It is thus difficult to attack (either via data poisoning or model poisoning) as the validation set of each client is private.
@@ 4.4 Asset Redistribution
As there are two independent actions, there are two parallel reward-and-slash designs for proposing and voting. For both actions, the randomly selected proposers and voters are required to stake a fixed sum of tokens before they act. If some of these actors fail to stake (they do not have enough tokens left), they lose their access to the blockchain and are removed from the FL system permanently. Proposers will be rewarded with tokens accumulated in an independent pool (if there are any tokens left in the pool) if the global aggregation is approved and lose their stakes if the global aggregation is rejected. The reward-and-slash design for the proposers is illustrated in Algorithm 1. For the voters, the majority party will not only take back their stakes but also be rewarded with the staked tokens lost by the minority party. The reward-and-slash design for the voters is illustrated in Algorithm 2. In the following section, Sec. 4.5, we demonstrate that under the proposed design and assumptions in Sec. 3.3, malicious voters have no incentive to make dishonest votes.
@@
6:${pool}_{p}$: Pool for proposers
@@
7:${pool}_{v}$: Pool for voters

### v2

# Defending Against Poisoning Attacks in Federated Learning with Blockchain

@@ 1 Introduction
Though integrating blockchain with existing FL systems can partially solve the threat to the central server, it cannot guarantee the quality of uploaded model updates from the clients. That is to say, blockchain-enabled FL systems are still vulnerable to client-side malicious attacks [14]. In this work, we define malicious behaviors as actions that intentionally decrease the learning performance (e.g. missing accuracy and convergence) of the global model via poisoning attacks (such as data poisoning [15] or model poisoning [14]). Instead of hacking the central server, the attackers can sabotage the FL systems by manipulating the clients. This work focuses on defending against client-side poisoning attacks. One solution is to combine blockchain-enabled FL with cryptographic protocols, such as Fully Homomorphic Encryption (FHE) [16] and Secure Multi-Party Computation (SMPC) [17], to mitigate malicious behaviors from the client side. However, the adoption of these intricate cryptographic protocols introduces significant computational overhead for FL participants, thus impairing the system performance. Besides, the malicious clients can still attack the system without breaching the protocols. It is challenging to address malicious behaviors without substantially compromising the efficiency of a blockchain-based FL system.
@@
We propose a generic framework that can integrate an FL system with a blockchain system and can defend against poisoning attacks without adopting complex cryptographic protocols. The proposed defense mechanism is motivated by proof-of-stake (PoS) [18], a consensus mechanism in blockchain, and The Resistance [19], a role-playing board game. PoS has an incentive mechanism that encourages honest behaviors by rewarding it and punishes dishonest behaviors via slashing. The Resistance, on the other hand, has two mismatched competing parties, where the party with a larger size is denoted as the resistance force and the other party is denoted as the spies. In The Resistance, there is a voting mechanism where, in each round, each player conducts independent reasoning and votes for a player, and the player with the highest votes will be deemed as a “spy” and kicked out of the game. The goal of the resistance force is to vote out all the spies while the spies aim to impersonate the resistance force and survive until the end. Based on these two concepts, this work proposes a novel majority-voting mechanism for global aggregation where each participating client independently validates the quality of aggregated local updates and votes for acceptance of the global update. The aggregation mechanism is stake-based where participating clients stake assets [^1] In practice, the staked assets can be linked with cryptocurrency or real currency to increase the financial cost of malicious attacks. or tokens (a quantitative measurement of the asset, which can be used to indicate the trustworthiness of the client in our system) for their own actions. There are two types of actions, proposing (uploading local updates) and voting. If the majority vote is to accept the global aggregation, a proposer will be refunded with its staked tokens and a voter who votes for acceptance will not only be refunded but also be rewarded with the staked tokens from the voters who vote for rejection, and vice versa. The overall procedure of the stake-based aggregation mechanism is illustrated in Fig. 1. To the best of our knowledge, this is the first work that integrates the majority voting and incentive mechanisms in the FL and blockchain literature.
@@
We evaluate the proposed framework on a practical financial problem, namely loan default prediction. We simulate the FL and blockchain environment for the Lending Club Kaggle challenge dataset and ChestX-ray14 dataset [20], to conduct experiments in a controllable setting and to provide insights into the problem of interest. We empirically show that an FL system can maintain robust performance under malicious attacks by introducing the proposed stake-based aggregation mechanism.
@@
We formulate the problem of decentralized federated learning with blockchain in the presence of poisoning attacks.
For the first time, we introduce a novel stake-based aggregation mechanism designed to fortify federated learning systems against poisoning attacks. In contrast to prior solutions, our mechanism boasts the distinct advantage of seamless integration into any blockchain enabled with smart contracts, all without necessitating alterations to the foundational consensus structure of the underlying blockchain. This approach not only enhances security but also simplifies accessibility, rendering it a more user-friendly option for federated learning participants.
We evaluate the robustness of the proposed framework in a simulated environment and provide initial empirical insights into the problem of interest. The findings show evidence that stake-based FL is an under-explored research problem with potential advantages compared with existing FL paradigms in terms of defending against poisoning attacks.
The rest of the paper is organized as follows. Sec. 2 reviews the related work on FL and blockchain computing. Sec. 3 formulates the problem of interest, defines the key concepts, and lists the necessary assumptions. Sec. 4 presents the method and theoretical result. Sec. 5 and Sec. 6 provide the experimental details on two different setups. Sec. 7 concludes this work.
Table 1: Comparison of existing blockchain-based FL solutions.
Framework Technique Incentive Mechanism Cryptocurrency Incentive Blockchain Agnostic Attack Mitigation
[16] FHE -
[17] SMPC -
[21] DP -
[22] DP-GAN -
[23] incentive mechanism -
[24] reputation + reverse auction -
[25] blockchain + reward only
[13] committee consensus
[26] committee consensus
Ours majority vote + reward & slash
@@ 2 Related Work
In this section, we review the recent progress on blockchain-based FL, and highlight the difference between the proposed method and existing studies.
@@ 2.1 Blockchain
Blockchains refer to distributed ledgers that operate on a global peer-to-peer (P2P) network, as exemplified by popular cryptocurrencies such as Bitcoin [27] and Ethereum [9]. Users can freely join or leave the blockchain system, without a central authority in place to ensure common agreement on the distributed ledgers. Instead, users rely on consensus protocols [18, 28], such as proof-of-work (PoW) or PoS, to achieve agreement in a distributed setting.
As shown in Fig. 2, a blockchain transaction typically involves a sender who transfers digital assets, such as cryptocurrencies, to a recipient. The sender authorizes the transaction with a digital signature combining transaction details and their private key. The transaction is then broadcasted over a P2P network to miners, who are participants in the network responsible for verifying and adding new blocks of transactions to the blockchain. Miners validate and confirm the transaction using consensus protocols, to ensure that the transaction is legitimate and not a duplicate or fraudulent transaction. Once confirmed, the transaction is added to a block, which is then linked to the previous block using cryptographically hash functions [29], forming a chain of blocks (i.e. missing, blockchain). The block is then propagated to all the participants in the network, creating a decentralized, immutable record of the transaction. The combination of cryptography and consensus protocols enhances the security, transparency, and decentralization of transactions, underscoring blockchain’s potential across various applications [30, 11, 12, 31, 32].
Another key feature of blockchain technology is the use of smart contracts [9], which are quasi-Turing-complete programs that can be executed within a virtual machine. When a transaction is initiated, a smart contract is typically used to encode the terms and conditions of the transaction, such as the amount, currency, and time of transfer. The smart contract is then stored on the blockchain network and executed automatically when the predefined conditions are met.
@@ 2.2 Federated Learning with Blockchain
Traditional FL faces challenges [33, 34, 35], such as privacy and security concerns (e.g. missing, poisoning attacks), unreliable communication, and difficulty in reaching a consensus among the parties. Blockchain, on the other hand, provides a decentralized, secure, and transparent platform for data storage and sharing. This makes the use of blockchain for FL a promising direction to potentially address privacy and security concerns by allowing parties to keep their data private while still contributing to the training process. Additionally, blockchain can provide a secure communication channel for FL participants and ensure the integrity of the FL process.
Current blockchain-based FL designs [35, 36, 8, 37, 38] have been broadly used in diverse fields [39, 40, 41]. For example, Ma et al. [25] propose a blockchain-assisted decentralized FL (BLADE-FL) framework, to prevent malicious clients from poisoning the learning process. Li et al. [13] analyze the impact of lazy clients on the learning performance of BLADE-FL and propose optimization for minimizing the loss function. Cui et al. [22] propose a blockchain-based decentralized and asynchronous FL framework for anomaly detection in IoTs by using a model named DP-GAN. Qu et al. [26] introduce a committee-based blockchain consensus algorithm for decentralized FL to prevent the single point of failure and poisoning attacks in FL.
Despite the potential benefits of combining FL with blockchain, several challenges remain. For instance, FL systems are still vulnerable to client-side malicious attacks [14] and lack incentive-compatible mechanisms to motivate FL participants to behave honestly during the training process. For instance, Miao et al. [16] leverage cosine similarity and blockchain to counteract poisoning attacks and penalize malicious clients. Their designs also rely on cryptographic techniques such as FHE. Kalapaaking et al. [17] adopt SMPC to enhance the security of blockchain-based FL which can mitigate poisoning attacks for healthcare systems. Yan et al. [21] integrate differential privacy (DP) techniques with blockchain to mitigate the issues of a single point of failure or untrusted aggregation caused by a malicious central server in privacy-preserving federated learning. However, using FHE, SMPC, or DP, instead of incorporating incentive mechanisms, may introduce an additional computation burden for clients or the server.
Several incentive mechanisms [23, 24] have recently been proposed to encourage participants and enhance model accuracy in blockchain-based FL. However, it remains unclear how to effectively utilize the blockchain infrastructure and leverage its inherent incentive mechanism (i.e. missing cryptocurrencies) to incentivize trustworthy FL behaviors and penalize malicious clients. Furthermore, to thwart potential malicious activities, such as poisoning attacks, existing blockchain-based FL solutions [8, 36, 25, 13, 26] have pointed out that participants can engage in both the FL training process and the block validation in PoS-based blockchains or mining activities in PoW-based blockchains. These design choices not only raise the entry barrier for regular users wishing to partake in blockchain-based FL systems but also add complexity to the fundamental consensus mechanism.
In this work, we introduce a PoS-based reward-and-slash mechanism for the FL system. We compare our solution with existing work in Table 1. Our solution can be seamlessly integrated into any smart-contract-enabled blockchain without requiring modifications to the underlying consensus design (i.e. missing, our design is blockchain agnostic). This approach facilitates the participation of any FL users, making it more accessible and user-friendly.
@@ 3 Problem Formulation
This section introduces the problem of interest, the definition of the malicious behaviors considered, and the underlying assumptions in this work. The main definitions and notations adopted in this work are summarized in Table 2.
@@ 3.3 Assumptions
A4: There is no capacity constraint on the hardware, including computing, communication, and storage, allowing us to solely focus on the algorithmic side of the problem.
A5: The underlying blockchain of the FL system of interest is running securely with a consensus protocol that ensures the validity and integrity of transactions and blocks. While the security of the blockchain is crucial for the overall security of the FL system, addressing the malicious miners falls outside the scope of this study.
Table 2: Summary of main definitions and notations.
Notation Description
blockchain-based FL FL framework with a blockchain architecture
FedAVG Federated Averaging [6], a FL algorithm
non-IID Non-independently and identically distributed
$\mathcal{K}$ Set of participating clients at round $t$
$K_{v}^{t}$ The selected set of voters
$\mathcal{K}_{p}^{t}$ Set of proposers at round $t$
$a^{t}$ Majority voting decision at round $t$
$M_{k}$ Asset of client $k$
$\gamma_{p}$ Staked tokens for proposing
${pool}_{p}$ Pool for storing proposers’ stake
@@ 4 Method
In this section, we first introduce the basics of federated aggregation in Sec. 4.1, and describe the local validation and majority voting in Sec. 4.2 and Sec. 4.3. We then propose a novel incentive mechanism in Sec. 4.4. A theoretical analysis on malicious voting is presented in Sec. 4.5. We then describe the whole training pipeline in Sec. 4.6. Finally, the analysis of computational cost is provided in Sec. 4.7.
Figure 3: A round-based training process. In the initial state (indexed as ①), both honest (black) and malicious (red) clients exist in an FL system. In the final state (indexed as ⑥), all malicious clients are expected to be removed from the system. To reach the final state from the initial state, multiple rounds of training are required. Here are two possible scenarios, the proposed aggregation is either approved (the upper branch) or denied (the lower branch) by the voters. In each round (within the dotted blue line), a subset of clients are randomly selected as proposers, and another subset of clients are randomly selected as voters. The proposers and voters interact with the blockchain following the order of orange arrows (from ② to ⑤).
@@ 4.2 Local Validation
In contrast to standard FL algorithms, the aggregated global model is not recorded in a block directly. Instead, $\tilde{\theta}_{0}^{t}$, a copy of $\theta_{0}^{t}$ is downloaded by a randomly selected set of clients, denoted as voters, $\mathcal{K}_{v}^{t}$. A voter $k$ runs a local inference with $\tilde{\theta}_{0}^{t}$ on its local test set and outputs a local validation score. The local validation score $s^{t}_{k}$ is a scalar, which can be linked with common metrics of ML tasks [^2] For example, common evaluation metrics include accuracy for classification, mean Intersection over Union (mIOU) for semantic segmentation, and mean average precision (mAP) for object detection.. If $s^{t}_{k}$ is not lower than a threshold, the voter votes for accepting this aggregated model; otherwise, the voter votes against it. The threshold can be based on a validation score $s^{t-1}_{k}$ acquired in the previous round. In the training of ML tasks, the scores can be volatile due to the characteristics of the tasks. Thus, a hyperparameter $\epsilon\in(0,1)$ is introduced to control the tolerance of performance decrease in a single round. Mathematically, the $k$-th voter has the following score.
@@
It is worth mentioning that the likelihood of the attackers consistently manipulating the scores by fooling all the randomly selected voters (e.g. missing via adversarial attacks [45]) diminishes quickly towards zero as the number of epochs increases. According to A4, the majority of voters are honest. It is thus difficult to attack (either via data poisoning or model poisoning) as the validation set of each client is private.
@@ 4.4 Asset Redistribution
As there are two independent actions, there are two parallel reward-and-slash designs for proposing and voting. For both actions, the randomly selected proposers and voters are required to stake a fixed sum of tokens before they act. If some of these actors fail to stake (they do not have enough tokens left), they lose their access to the blockchain and are removed from the FL system permanently. Proposers will be rewarded with tokens accumulated in an independent pool (if there are any tokens left in the pool) if the global aggregation is approved and lose their stakes if the global aggregation is rejected. The reward-and-slash design for the proposers is illustrated in Algorithm 1. For the voters, the majority party will not only take back their stakes but also be rewarded with the staked tokens lost by the minority party. The reward-and-slash design for the voters is illustrated in Algorithm 2. In the following section, Sec. 4.5, we demonstrate that under the proposed design and assumptions in Sec. 3.3, malicious voters have no incentive to make dishonest votes. Note, Sec. 4.4 highlights the key difference between the proposed voting mechanism and traditional majority voting as the voting is directly linked with the incentive mechanism.
@@
6:${pool}_{p}$: Pool for storing proposers’ stake
@@
7:${pool}_{v}$: Pool for storing voters’ stake